OpenStudy (anonymous):

simplify (cos u + sin u)2

OpenStudy (vishweshshrimali5):

First of all apply this identity: \[\large{(a+b)^2 = a^2 + b^2 + 2ab}\] Here, a = cos u and b = sin u

OpenStudy (vishweshshrimali5):

Can you do this part ?

OpenStudy (vishweshshrimali5):

Now, you will also need this identity: \[\large{\sin^2{\theta} + \cos^2{\theta} = 1}\]

OpenStudy (vishweshshrimali5):

Can you solve it now @mlgossett ?

OpenStudy (perl):